# thanks: Haihao Shi and Zhenyang Huang contribute equally to this workthanks: Corresponding authorthanks: Corresponding author

### v1

@@ VII Discussion
Our investigation reveals a compelling trade-off between predictive accuracy and model interpretability in describing post-FFC neutrino dynamics. As shown in Table 2, the parametric KAN sets a new benchmark for accuracy, surpassing previous MLP-based models in total errors. However, this peak performance comes at the cost of interpretability. The process of deriving an explicit symbolic formula from the KAN inevitably introduces inaccuracies. We identify three primary sources for this precision loss: (i) the intrinsic difficulty in fitting certain learned B-spline activation functions to elementary functions with high fidelity; (ii) our strategy of averaging B-spline parameters across multiple runs to ensure robustness , which means the final model does not represent the single best-performing run; and (iii) the fact that even with averaging, some activation functions did not achieve a coefficient of determination ($R^{2}$) greater than 0.99 during symbolic regression.
@@
On the other end of the spectrum, PySINDy delivers what the symbolic KAN struggles with: a remarkably concise, low-rank model in the form of simple governing equations. This elegance and simplicity, however, are paid for with lower predictive accuracy compared to the neural network frameworks. We hypothesize this accuracy gap stems from the differing flexibility of the optimization frameworks. Modern neural network architectures, like our KAN, can seamlessly integrate complex, physics-based prior knowledge. For instance, our KAN model successfully incorporated three distinct physics-inspired loss terms via a dynamic weighting mechanism to enforce physical conservation laws. Such intricate, multi-term constraints are not as easily implemented within the standard PySINDy workflow, potentially limiting its ability to converge to a more physically precise solution.
Ultimately, our work demonstrates that there is no single “best” interpretable model for this problem; rather, the optimal choice depends on the scientific objective. For developing high-fidelity surrogate models intended for integration into larger hydrodynamic simulations where precision is paramount, the parametric KAN or MLP is the superior choice. Conversely, for the goal of gaining physical insight, generating new hypotheses, and discovering the simplest possible governing laws of the system, PySINDy is unparalleled, despite its reduced accuracy. This comparative analysis provides a valuable roadmap for researchers applying machine learning in astrophysics, clarifying the distinct advantages of different interpretable techniques. Future work could focus on developing hybrid methods that leverage the strengths of both approaches. For instance, one could use the sparse governing equations from SINDy as a regularization term, or incorporate insights from GSA to guide the learning of a KAN, thereby pursuing the dual goals of accuracy and interpretability simultaneously.
@@ VIII Conclusion
In this work, we use and compare two interpretable frameworks, KANs and SINDy, as alternatives to ’black-box’ models like MLPs for FFCs in supernova neutrinos. Based on these models, we predict post-FFC spectra using only the first two radial moments from modern CCSNe and NSM simulations, which are typically available in large-scale simulations. This provides a pathway to embed FFC physics into these simulations via surrogate models that are not only accurate but also more physically transparent. Our results suggest that, for such complex physical systems, there can be a trade-off between prediction accuracy and model interpretability.
Our analysis shows that the parametric KAN offers the highest prediction accuracy and serves as an excellent alternative to standard MLP models. However, when expressed in symbolic form through regression, the KAN model sacrifices some accuracy and becomes more complex, which can limit its direct use for physical insight. In contrast, PySINDy performs well where symbolic KAN struggles. It produces a simpler and more elegant low-rank dynamical model, though this simplicity comes at the cost of lower predictive accuracy compared to neural network methods. Despite these trade-offs, both interpretable frameworks successfully identify the same key physical feature: the relatively weak dependence on energy as well as on $F_{\nu_{e}}^{\mathrm{tot}}$ and $F_{\bar{\nu}_{e}}^{\mathrm{tot}}$ in the post-FFC neutrino first two radial moments. Moreover, PySINDy results highlight that the number density of heavy-lepton neutrinos is the main factor influencing FFC evolution. These interpretable findings are consistent with existing theoretical predictions and underscore the value of moving beyond black-box models.
In summary, our research enhances the effectiveness of NN in predicting the asymptotic outcomes of FFCs in a multienergy neutrino gas. Leveraging this identified structure, the post-conversion neutrino energy spectra are reconstructed with an average accuracy of up to 90%. Furthermore, we employ interpretable models to analyze key factors influencing the evolution during FFCs. Future work should focus on extending the study to more realistic neutrino gases with non-axisymmetric angular distributions, where $\nu_{x}$ and $\bar{\nu}_{x}$ may exhibit distinct behaviors. Interestingly, recent work provides a systematic treatment of inhomogeneous many-body (MB) dynamics of the fast flavor instability, demonstrating that MB effects are expected to persist even in thermodynamically large neutrino ensembles [87]. These advancements will improve the feasibility of incorporating FFCs into CCSNe and NSM simulations, thereby enhancing our ability to accurately model and predict these extreme astrophysical phenomena.
@@ Acknowledgements.
Thanks to Sajad Abbar for providing the dataset used in this study and for his guidance, as well as to Shun Zhou and Xuwei Zhang for their useful discussions. This work received support from the National Natural Science Foundation of China under grants 12288102, 12373038, 12125303, 12090040/3, U2031204 and 12433009 (with Junda Zhou as the sole recipient); the Natural Science Foundation of Xinjiang No. 2022TSYCLJ0006; the science research grants from the China Manned Space Project No. CMS-CSST-2021-A10; the National Key R&D Program of China No. 2021YFA1600401 and No. 2021YFA1600403; the Natural Science Foundation of Yunnan Province Nos. 202201BC070003 and 202001AW070007; the International Centre of Supernovae, Yunnan Key Laboratory No. 202302AN360001; and the Yunnan Revitalization Talent Support Program -- Science & Technology Champion Project No. 202305AB350003.
@@
The source code for the methods implemented in this study will be made publicly available in a public repository upon acceptance of the manuscript.

### v2

@@ VII Discussion
Our investigation reveals a notable trade-off between predictive accuracy and model interpretability in describing post-FFC neutrino dynamics. As shown in Table 2, the parametric KAN sets a new benchmark for accuracy, surpassing previous MLP-based models in total errors. However, this peak performance comes at the cost of interpretability. The process of deriving an explicit symbolic formula from the KAN inevitably introduces inaccuracies. We identify three primary sources for this precision loss: (i) the intrinsic difficulty in fitting certain learned B-spline activation functions to elementary functions with high fidelity; (ii) our strategy of averaging B-spline parameters across multiple runs to ensure robustness , which means the final model does not represent the single best-performing run; and (iii) the fact that even with averaging, some activation functions did not achieve a coefficient of determination ($R^{2}$) greater than 0.99 during symbolic regression.
@@
On the other end of the spectrum, SINDy delivers what the symbolic KAN struggles with: a remarkably concise, low-rank model in the form of simple governing equations. This elegance and simplicity, however, are paid for with lower predictive accuracy compared to the neural network frameworks. We hypothesize this accuracy gap stems from the differing flexibility of the optimization frameworks. Modern neural network architectures, like our KAN, can seamlessly integrate complex, physics-based prior knowledge. For instance, our KAN model successfully incorporated three distinct physics-inspired loss terms via a dynamic weighting mechanism to enforce physical conservation laws. Such intricate, multi-term constraints are not as easily implemented within the standard SINDy workflow, potentially limiting its ability to converge to a more physically precise solution.
Ultimately, our work demonstrates that there is no single “best” interpretable model for this problem; rather, the optimal choice depends on the scientific objective. For developing high-fidelity surrogate models intended for integration into larger hydrodynamic simulations where precision is paramount, the parametric KAN or MLP is the superior choice. Conversely, for the goal of gaining physical insight, generating new hypotheses, and discovering the simplest possible governing laws of the system, SINDy is unparalleled, despite its reduced accuracy. This comparative analysis provides a valuable roadmap for researchers applying machine learning in astrophysics, clarifying the distinct advantages of different interpretable techniques. Future work could focus on developing hybrid methods that leverage the strengths of both approaches. For instance, one could use the sparse governing equations from SINDy as a regularization term, or incorporate insights from GSA to guide the learning of a KAN, thereby pursuing the dual goals of accuracy and interpretability simultaneously.
@@ VIII Conclusion
In this work, we use and compare two interpretable frameworks, KANs and SINDy, as alternatives to ’black-box’ models like MLPs for FFCs in supernova neutrinos. Based on these models, we predict post-FFC spectra using only the first two radial moments from modern CCSNe and NSM simulations, which are typically available in large-scale simulations. This provides a pathway to embed FFC physics into these simulations via surrogate models that are not only accurate but also more physically transparent. Our results suggest that, for such complex physical systems, there can be a trade-off between prediction accuracy and model interpretability. Despite these trade-offs, both interpretable frameworks successfully identify the same key physical feature: the relatively weak dependence on energy as well as on $F_{\nu_{e}}^{\mathrm{tot}}$ and $F_{\bar{\nu}_{e}}^{\mathrm{tot}}$ in the post-FFC neutrino first two radial moments. Moreover, the derivatives-based contribution heatmaps of KAN and SINDy results highlight that the initial number density of heavy-lepton neutrinos is the main factor influencing FFCs evolution results. These interpretable findings demonstrate the benefits of interpretability and underscore the value of moving beyond black-box models.
In summary, our research enhances the effectiveness of NN in predicting the asymptotic outcomes of FFCs in a multienergy neutrino gas. Leveraging this identified structure, the post-conversion neutrino energy spectra are reconstructed with an average accuracy of up to 90%. Furthermore, we employ interpretable models to analyze key factors influencing the evolution during FFCs. Future work should focus on extending the study to more realistic neutrino gases with richer physical details. Interestingly, recent work provides a systematic treatment of inhomogeneous many-body (MB) dynamics of the fast flavor instability, demonstrating that MB effects are expected to persist even in thermodynamically large neutrino ensembles [85]. These advancements will improve the feasibility of incorporating FFCs into CCSNe and NSM simulations, thereby enhancing our ability to accurately model and predict these extreme astrophysical phenomena.
@@ Acknowledgements.
Thanks to Sajad Abbar for providing the dataset used in this study and for his guidance, as well as to Kaifan Ji, Shun Zhou and Xuwei Zhang for their useful discussions. This work received support from the National Natural Science Foundation of China under grants 12288102, 12373038, 12125303, 12090040/3, U2031204 and 12433009 (with Junda Zhou as the sole recipient); the Natural Science Foundation of Xinjiang No. 2022TSYCLJ0006; the science research grants from the China Manned Space Project No. CMS-CSST-2021-A10; the National Key R&D Program of China No. 2021YFA1600401 and No. 2021YFA1600403; the Natural Science Foundation of Yunnan Province Nos. 202201BC070003 and 202001AW070007; the International Centre of Supernovae, Yunnan Key Laboratory No. 202302AN360001; and the Yunnan Revitalization Talent Support Program - Science & Technology Champion Project No. 202305AB350003.
@@
The source code for the interpretable data-driven methods presented in this study, including the implementation of our KAN and SINDy frameworks, is available in an anonymous repository  [^1] https://anonymous.4open.science/r/interpretable-data-driven-methods-for-FFCs-1827/.